Part 1. Choose the correct answer and write the letter responding to it on the answer sheet.


1. The type of graph for grouped frequency distribution that plot marks against frequency

A. O gives                           B. Vertical line graph

C.Frequency polygon        D. Histogram

E. None


2 In a histogram, the bar with the greatest height represents

A. The median group

B. Interval containing the mean value

C.The modal group

D. Interval with greatest variation

E. None


3. If the price of a commodity increasca from birr 16 to birr 81 in a period of four years

Then the average percentage increase in price per year is

A. 50%          B. 65%          C. 16.25%       D.50.6%


4.The harmonic mean of the numbers 1, 2, 3,6,12 is

A. 4.8         B. 24          C. 10          D. 3.6


5. The mean age of students in three sections consisting of 38, 30. and 32 students are 35, 29, and 25 years respectively then mean age of all the students taken together is

A. 33.3 yr      B. 29.7 yr       C. 29.5 yr      D. 30.0 yr





1
Expert's answer
2022-02-03T11:54:16-0500

Solution:

  1. C.Frequency polygon
  2. C.The modal group
  3. the average percentage increase in price per year"=\dfrac{81-16}{16\times 4}\times 100\%=101.5625\%"

No correct option.

4. The harmonic mean of the numbers 1, 2, 3,6,12 "=\dfrac{n}{\frac{1}{a_1}+\frac{1}{a_2}+...+\frac{1}{a_n}}"

"=\dfrac{5}{\frac{1}{1}+\frac{1}{2}+\frac{1}{3}+\frac{1}{6}+\frac{1}{12}}
\\=\dfrac{5}{\frac{25}{12}}
\\=\dfrac{12}{5}=2.4"

B is correct. It is written incorrectly.

5. mean age of all the students taken together is "\dfrac{38\times35 + 30\times29 + 32\times 25}{38+30+32}=30"

D is correct.